OpenStudy (anonymous):

drivers looking for a sunken ship have defines the search area as a tringle with adjacent sides of length 2.75 miles and 1.32 miles. the angle between the sides of the triangle is 35degrees. to the nearest hundredth, find the search area...medal,

OpenStudy (anonymous):



OpenStudy (anonymous):

using law of cosines (x^2 = y^2 + z^2 - 2yzcos(angle)), you get that x = 1.832m

OpenStudy (anonymous):

here they probably want you to find the height to find the area. There is something called herons formula. it says Area of a triangle = sqrt(p(p-a)(p-b)(p-c)) were p is half the perimeter. so p = 2.951m. Now area of triangle = 1.04miles